OpenStudy (anonymous):

What is the area of a sector with radius equal to 8 and measure of arc equal to 300°? HELP ME

OpenStudy (anonymous):

300° = 300*(3.14/180) radians = 5.233 radians area of sector = 1/2 * r^2 * angle subtended = 1/2 * 64 * 5.233 = 167.47

OpenStudy (anonymous):

angle 300 deg 22 30 * 22 * 64 42240 area = -------- * πr² = ---------- * ----- * (8)² = ------------ = ------- 360 deg 360 deg 7 36 * 7 252 area = 167.62 cm²